OpenStudy (luckycoins888):

FAN + MEDAAALLL!!!!!!!! The surface area of a cube is 486 ft2. What is the area of one face of this cube? ft2

OpenStudy (danjs):

A cube has 6 faces right?

OpenStudy (danjs):

All of the faces are the same size, like on a pair of dice

OpenStudy (danjs):

Each face is a square with area Length x width

OpenStudy (danjs):

The total area of the entire cube is 6 of those squares Surface Area Cube = 6 x Length x Width

OpenStudy (danjs):

A squares Length is the same as the Width, so you can write Surface Area Cube = 6 x (Side Length)^2

OpenStudy (danjs):

486 = 6*(side Length)^2 81 = (Side Length )^2 Side Length = 9 One Side Area = 1 x (Side)^2 One Side Area = 9^2 = 81 ft^2

OpenStudy (danjs):

Or you could quickly solve: Total SA = 486 One Side is 1/6 of the total surface area, One side has area 486/6 = 81 ft^2

OpenStudy (danjs):

everything clear?

OpenStudy (luckycoins888):

yep

OpenStudy (anonymous):

was dis one answerd already?